OpenStudy (anonymous):

medal An airplane is cruising at a speed of 325 mph. Which direct variation equation represents this situation? (Note: The variable d represents distance, and t represents time.)

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (tkhunny):

Usually, Distance = Rate * Time

OpenStudy (anonymous):

d=1/325 d=325t d=t/325 d=t-325

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@KyanTheDoodle

OpenStudy (anonymous):

i think its d=t/325

OpenStudy (anonymous):

@vera_ewing

OpenStudy (tkhunny):

Which one looks like what I wrote?

OpenStudy (anonymous):

1/325 t

OpenStudy (tkhunny):

Mine has only multiplication.